Eugene's population was 145,000 in 2005. Find the percent growth from 2005 to 2015.
Answer by fractalier(6550)

Assuming that there are 161000 people in 2015, as shown in the previous problem, the percent growth is found by
% change = (new - old)/old = (161000-145000)/145000 = 11% increase in ten years